OpenStudy (anonymous):

4m + -9 = -17 How do I solve?

OpenStudy (davidusa):

4n-9=-17 add 9 to both sides. 4n =-17+9

OpenStudy (anonymous):

These types of problems you get an idea what you need to do So for your equation 4m + -9 = -17 You would add 9 to each side making your answer look like this : m=-2 Hope this helped!

OpenStudy (mathstudent55):

@holly01 Aren't you skipping a few steps from adding 9 to both sides to 4m + -9 = -17 to arriving at m = -2? What do you get when all you do is add 9 to both sides?

OpenStudy (anonymous):

I am very sorry So the first step you take is of course take 9 from each side looking like this 1. 17+9 = 26 So your answer is completely wrong But like @DavidUsa said you have 4n = 17+9